Given two bases



B={1−x,2+x,3−x+x2}



and



C={1,2+x,1+x−x2}



of P2, the vector space of polynomials of degree ≤2,



(i) find p(x)∈P2 whose coordinates with respect to B is [p(x)]B=⎡⎣⎢1 −1 3⎤⎦⎥,



(ii) find the transition (change of coordinates) matrix CMB∈R3×3 from B to C,



(iii) calculate the coordinates [p(x)]C∈R3 of p(x)∈P2 with respect to C.


1
Expert's answer
2022-01-02T16:49:50-0500

Solution:

(i):

The coordinates of p(x)"\in" P2 with respect to the basis B is "[p(x)]_B=\begin{bmatrix} 1 \\-1\\3\end{bmatrix}"

"\therefore p(x)=1(1-x)-1(2+x)+3(3-x+x^2)
\\=(1-2+9).1+(-1-1-3)x+3.x^2
\\=8-5x+3x^2"

(ii):

Let "1-x=a.1+b(2+x)+c(1+x-x^2)"

"\Rightarrow 1-x=(a+2b+c)+(b+c)x-cx^2"

On comparing, we get,

"a+2b+c=1
\\b+c=-1
\\-c=0
\\\Rightarrow c=0,b=-1,a=3"

So, "1-x=3(1)-1(2+x)+0.(1+x-x^2)\ ...(i)"

Also, "2+x=0(1)+1(2+x)+0.(1+x-x^2)\ ...(ii)"

Also, "3-x+x^2=4(1)+0(2+x)+(-1).(1+x-x^2)\ ...(iii)"

From (i), (ii), (iii), we get the transition matrix from B to C as:

"_CM_B=\begin{bmatrix} 3&0&4 \\-1&1&0\\0&0&-1\end{bmatrix}"

(iii):

From part (i), we have "p(x)=8-5x+3x^2"

Let "8-5x+3x^2=p(1)+q(2+x)+r(1+x-x^2)"

"\Rightarrow 8-5x+3x^2=(p+2q+r)+(q+r)x-rx^2"

On comparing, we get,

"p+2q+r=8
\\q+r=-5
\\-r=3
\\\Rightarrow r=-3,q=-2,p=15"

So, we get, "p(x)=15(1)-2(2+x)-3(1+x-x^2)"

So, the coordinates of p(x) w.r.t basis C is:

"[p(x)]_C=\begin{bmatrix} 15 \\-2\\-3\end{bmatrix}"